Pops Cycle Shop sells bicycles and tricycles. The number of bicycles is 1 less than 4 times the number of tricycles. All the bicycles and tricycles together have a total of 174 wheels. How many bicycles are there?
Answer by Alan3354(21608)

Pops Cycle Shop sells bicycles and tricycles. The number of bicycles is 1 less than 4 times the number of tricycles. All the bicycles and tricycles together have a total of 174 wheels. How many bicycles are there?
----------------------
b = # of bicycles
t = # of tricycles
--------------
b = 4t - 1 (The number of bicycles is 1 less than 4 times the number of tricycles.)
2b + 3t = 174 (total wheels, 2 per bike, 3 per tricycle)
Sub for b into 2nd eqn
2(4t-1) + 3t = 174
8t-2 + 3t = 174
11t = 176
t = 16
------------
b = 4t-1 = 4*16-1
b = 63